# Multi-target Range, Doppler and Angle estimation in MIMO-FMCW Radar with Limited Measurements

### v1

@@ I Introduction
Sensing the environment using radar or other sensors is an integral part of many engineering applications. A radar usually localizes a target of interest by estimating its range, Doppler velocity, and angle of arrival (AoAs). Owing to their portability, low cost, and high range resolution, frequency-modulated continuous wave (FMCW) radars are often preferred in short-range applications like advanced driving assistance systems (ADAS)[2, 3], synthetic aperture radars (SARs)[4, 5], surveillance systems[6, 7], and human vital sign monitoring[8]. FMCW radars transmit a finite number of linear frequency-modulated (LFM) chirps in each coherent processing interval (CPI). The receiver mixes the signal reflected from targets with the transmitted signal, producing a complex sinusoidal intermediate frequency (IF) signal. The IF signal’s frequencies then determine the target ranges and velocities. To further localize targets in the angular domain, an array of multiple transmitter and receiver antennas is required. Multiple-input multiple-output (MIMO) radars transmit multiple orthogonal waveforms and jointly process the target returns across multiple receivers, providing additional degrees of freedom and higher resolution compared to conventional phased array radars, but with fewer physical antenna elements.
@@
I-A Prior Art
Earlier CS applications in radar systems typically focused on uniform sampling or full measurements. In [17], CS replaced the matched filter reducing the sampling rate for radar imaging, while [18, 19] explored high-resolution estimation in a general 2D time-frequency domain. Efficient waveforms to reconstruct sparse targets were suggested in [20]. In [21, 22], the randomized measurement framework was utilized for radar imaging in remote sensing and inverse SAR applications. A CS-based constant false alarm rate (CFAR) detector was developed in [23] while [24] used CS to leverage the diversity of various array orientations for high-resolution spatial spectrum reconstruction. Recently, [25] proposed a two-level CS-based compression technique for fast target detection in MIMO radar compared to conventional methods.
@@ I-A Prior Art
A key benefit of CS techniques is the reduction in measurement requirements for sparse signals. In radar, the target scene is sparse because only a few targets are typically present, making CS a natural fit for the problem. As a result, CS has been widely adopted to enable sub-Nyquist radars with reduced measurements [26, 27, 28]. In the spatial domain, sparse linear arrays (SLAs), with fewer antenna elements compared to uniform linear arrays (ULAs), have been introduced for both pulsed and continuous-wave radars [9, 29, 30]. Optimal sparse array design was investigated in [30], while [9] designed a non-uniform SLA and applied digital beamforming for AOA estimation after interpolating the missing measurements. Conversely, [29] suggested matrix completion methods to reconstruct the corresponding linear array. Although these methods involve sparse measurements, they complete the missing measurements before parameter estimation, resulting in a high computational load at the receiver.
@@
Alternatively, spatial CS enables direct parameter recovery from SLAs[31, 32]. For instance, [31] focused on pulsed-MIMO radar and estimated AOAs for a specific range-velocity bin using different CS recovery algorithms. Additionally, [33] incorporated velocity estimation but not range. In [32], velocities and angles were estimated using measurements from a small number of randomly positioned transmitters and receivers on a circular disc. Besides spatial compression, CS techniques have also been applied in radar systems for interference mitigation [34], spectrum sharing [35], joint radar-communication systems[36], and multi-target shadowing effect mitigation[37]. Furthermore, [38] discussed the advantages and challenges of applying CS in radar, including clutter cancellation. However, these earlier works primarily concentrated on pulsed-wave radars, leaving CS-based joint parameter estimation in FMCW radars largely unexplored.
@@ I-B Our contributions
Preliminary results of this work appeared in our conference publication [1], where we addressed range and angle estimation only, without any theoretical guarantees. In this work, we consider a MIMO-FMCW radar with a random SLA that transmits only a subset of chirps per CPI and present a joint Doppler-angle estimation framework. The reduced number of antenna elements and chirps correspond to spatial and slow-time CS, respectively. Our main contributions in this paper are:
1) Range estimation: Prior works on CS-based radars have mainly focused on pulsed-wave radars and/or recovering target angles and velocities within a specific range bin. In this paper, we first present a separable mixture model for the IF signal, applicable to both full and sparse measurements. We propose two methods for range estimation: (a) DFT-focusing followed by binary integration[10], and (b) range-orthogonal matching pursuit (range-OMP). Binary integration improves detection at low SNRs, while OMP provides higher resolution and reduced computational costs; see Remarks 2-3.
2) Joint Doppler-angle estimation: We jointly estimate target velocities and AOAs within each detected range bin using both vectorized and joint 2D CS techniques. Our methods deliver performance comparable to that of a full measurement system, with user-defined grids allowing for higher velocity and angular resolutions. Additionally, processing fewer measurements reduces computational complexity and enables the radar to explore multiple angular sectors simultaneously in a single CPI.
@@
Throughout the paper, we reserve boldface lowercase and uppercase letters for vectors (column vectors) and matrices, respectively, and $\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$ (or simply $a_{i_{1}\leq i\leq i_{2}}$) denotes a set of elements indexed by an integer $i$. The notations $[\mathbf{A}]_{:,i}$, $[\mathbf{A}]_{i,:}$ and $[\mathbf{A}]_{i,j}$ denote the $i$-th column, $i$-th row and $(i,j)$-th element of matrix $\mathbf{A}$, respectively. The transpose/ Hermitian/conjugate operation is $(\cdot)^{T/H/*}$, expectation/probability is $\mathbb{E}/\mathbb{P}[\cdot]$, the outer product is $\otimes$ and the complement of a set is $(\cdot)^{c}$. The $l_{2/1/0}$ norm of a vector is $\|\cdot\|_{2/1/0}$. The notations supp, diag, and vec denote support, diagonal matrix, and vectorization, respectively. Also, $\mathbf{I}_{n}$ and $\mathbf{0}$ denote a ‘$n\times n$’ identity matrix and an all-zero matrix, respectively. We represent the circular-normal and real-valued normal distributions as $\mathcal{CN}(\bm{\mu},\mathbf{Q})$ and $\mathcal{N}(\bm{\mu},\mathbf{Q})$ (with mean $\bm{\mu}$ and covariance matrix $\mathbf{Q}$) while $\mathcal{U}[a,b]/\mathcal{U}\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$ represents a uniform/ discrete-uniform distribution over interval $[a,b]$/ set $\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$.
@@ II System Model
The transmitters transmit orthogonal LFM chirps of carrier frequency $f_{c}$, chirp rate $\gamma$, and chirp duration $T_{c}$. The wavelength is $\lambda=c/f_{c}$, where $c$ represents the speed of light. For simplicity, we consider time-domain orthogonality, i.e., the transmitters transmit the same signal with relative time shifts. Hence, the received signal components corresponding to various transmitters are readily separated at different receivers. Alternatively, orthogonal waveforms for MIMO-FMCW radars have been designed using beat frequency and chirp rate offsets [39, 40]. In [41], difference co-chirps are proposed for high-accuracy and low complexity range-Doppler estimation in FMCW radars. We consider a CPI of duration $T_{p}=P_{max}T_{c}$ corresponding to $P_{max}$ chirps. The standard radar system with full (slow-time) measurements transmits for the entire CPI. In this work, the radar transmits only $P$ out of $P_{max}$ chirps ($P<P_{max}$). We denote the sparse chirp set by $\mathcal{P}_{s}=\{\zeta_{p}\}_{1\leq p\leq P}$ where $\zeta_{p}$ are randomly drawn distinct integers from $\{0,1,\ldots,P_{max}-1\}$ and distributed as $\mathcal{P}_{p}$. On the other hand, in the standard radar, the set $\{\zeta_{p}\}=\{0,1,\ldots,P_{max}-1\}$ and the chirps are transmitted in a specific angular sector of interest. In our sparse radar setup, the time corresponding to $\mathcal{P}_{s}^{c}$, i.e., the non-transmitted chirps, can be utilized in other angular sectors, as depicted in Fig. 1b. Throughout the paper, we consider the target’s parameter estimation in a single sector only. The received signals from other sectors can be trivially processed in the same manner.
@@
[…]
@@
where $a_{k}$ is the complex amplitude proportional to the $k$-th target’s radar cross-section (RCS) and $\tau^{k}_{n,m,p}$ is the total time-delay in the $k$-th target’s return. The delay $\tau^{k}_{n,m,p}$ consists of range delay $\tau^{R}_{k}$, Doppler delay $\tau^{D}_{p,k}$ and angular delay $\tau^{\theta}_{n,m,k}$ as
@@
×$\;\;\times\exp{\left(-j\pi\gamma(\tau^{k}_{n,m,p})^{2}\right)}\exp{\left(j2\pi f_{c}\tau^{k}_{n,m,p}\right)},$
@@
×$\times\underbrace{\exp{\left(-j\pi\gamma(\tau^{k}_{n,m,p})^{2}\right)}}_{\textrm{Term-II}}\underbrace{\exp{\left(j2\pi f_{c}\tau^{k}_{n,m,p}\right)}}_{\textrm{Term-III}}.$ (2)
We, henceforth, deal with only discrete-time signals and use $t$ to denote the discrete-time index with $10\leq t\leq N-$ 1 where $N=\lceil f_{s}T_{c}\rceil$. In a MIMO-FMCW radar with $N_{T}$ transmitters, $N_{R}$ receivers and $P$ chirps, we obtain ‘$N_{T}N_{R}\times P$’ sampled measurements $\{y_{n,m,p}[t]\}_{1\leq n\leq N_{T},1\leq m\leq N_{R},1\leq p\leq P}$.
@@
Term-I: For practical FMCW radars with narrow-band assumption, we have $\tau^{R}_{k}\gg\tau^{D}_{p,k}$ and $\tau^{R}_{k}\gg\tau^{\theta}_{n,m,k}$ for all targets such that $\gamma\tau^{k}_{n,m,p}/f_{s}\approx\gamma\tau^{R}_{k}/f_{s}$. This can also be verified numerically for a radar target with the parameter values provided in Table I. We define the normalized beat frequency (due to range) as $\Omega_{R}^{k}\doteq\gamma\tau^{R}_{k}/f_{s}$ and approximate Term-I as $\exp{(j2\pi\Omega_{R}^{k}t)}$. This term is further examined in Section III-B using DFT-based focusing or OMP-recovery methods to obtain the range estimates.
Term-II: Again, using $\tau^{R}_{k}\gg\tau^{D}_{p,k}+\tau^{\theta}_{n,m,k}$, we have $(\tau^{k}_{n,m,p})^{2}\approx(\tau^{R}_{k})^{2}$ [42]. Hence, Term-II becomes $\exp{(-j\pi\gamma)(\tau^{R}_{k})^{2}}$, a constant for each target independent of chirp and antenna indices. In the 3D-mixture model (3), we include this term in $\widetilde{a}_{k}$ and is not the focus of the parameter estimation problem addressed in this paper.
@@
×$\;\;\times\exp{\left(j2\pi\frac{A\sin(\theta_{k})}{2\lambda}(\alpha_{n}+\beta_{m})\right)}.$
@@
With these approximations and including the noise term $w_{n,m,p}[t]$ to represent the interference and noises present in the radar measurements, (2) simplifies to
@@
×$\;\;\times\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}+w_{n,m,p}[t],$ (3)
where $\widetilde{a}_{k}=a_{k}exp(j\pi\gamma(\tau^{R}_{k})^{2})exp(-j2\pi f_{c}\tau^{R}_{k})$. Here, (3) consists of a separable mixture of three different complex exponentials such that the target range, velocity, and AOA estimation problem is equivalent to 3D frequency estimation.
Remark 1.

### v2

@@ I Introduction
Sensing the environment using radar or other sensors is an integral part of many engineering applications. A radar usually localizes a target of interest by estimating its range, velocity, and angle of arrival (AoA). Owing to their portability, low cost, and high range resolution, frequency-modulated continuous wave (FMCW) radars are often preferred in short-range applications like advanced driving assistance systems (ADAS)[2, 3], synthetic aperture radars (SARs)[4, 5], surveillance systems[6, 7], and human vital sign monitoring[8]. FMCW radars transmit a finite number of linear frequency-modulated (LFM) chirps in each coherent processing interval (CPI). The receiver mixes the signal reflected from targets with the transmitted signal, producing a complex sinusoidal intermediate frequency (IF) signal. The IF signal’s frequencies then determine the target ranges and velocities. To further localize targets in the angular domain, an array of multiple transmitter and receiver antennas is required. Multiple-input multiple-output (MIMO) radars transmit multiple orthogonal waveforms and jointly process the target returns across multiple receivers, providing additional degrees of freedom and higher resolution compared to conventional phased array radars, but with fewer physical antenna elements.
@@
I-A Prior Arts
Earlier CS applications in radar systems primarily aimed to enhance resolution while still relying on uniform sampling or full measurement setups. For instance, [17] replaced the matched filter by CS using random projections to reconstruct radar images, thereby lowering ADC demands. Other works, such as [18, 19, 20], explored pseudo-random phase-coded transmit waveforms to improve resolution in delay-Doppler and range-angle domains. In [21], random frequency-hopping was employed for transmit waveform design in MIMO radar and SAR, while [22] utilized randomized waveforms for super-resolution imaging. These approaches also utilized randomly selected measurements across time, frequency, or spatial domains. In [23], a CS-based energy detector was applied to the recovered outputs, while [24] exploited spatial diversity through varying array orientations for high-resolution spectrum estimation, with randomness introduced by the dynamic sensor positioning. More recently, [25] proposed a GLRT-based detection method operating directly on compressed measurements acquired through random projections and selective sampling. However, these works acquire measurements at high uniform sampling rates, similar to conventional radars, and process only the selected or linearly projected measurements. While this can reduce computational load and improve resolution, it does not significantly reduce hardware complexity. In contrast, we adopt a fundamentally different approach by using a reduced number of physical antenna elements randomly distributed over the array aperture and randomly transmitting only a limited set of chirps in each CPI. Hence, we achieve high-resolution performance while substantially lowering both computational and hardware requirements. Note that our random chirp setup also differs from the difference co-chirps proposed in [26], where the sparse chirp set arises from difference co-array structures, such as co-prime and nested chirps, instead of being transmitted randomly.
@@ I-A Prior Arts
A key benefit of CS techniques is the reduction in measurement requirements for sparse signals. In radar, the target scene is sparse because only a few targets are typically present, making CS a natural fit for the problem. As a result, CS has been widely adopted to enable sub-Nyquist radars with reduced measurements [27, 28, 29]. In the spatial domain, sparse linear arrays (SLAs), with fewer antenna elements compared to uniform linear arrays (ULAs), have been introduced for both pulsed and continuous-wave radars [9, 30, 31]. Optimal sparse array design was investigated in [31], while [9] designed a non-uniform SLA and applied digital beamforming for AOA estimation after interpolating the missing measurements. Conversely, [30] suggested matrix completion methods to reconstruct the corresponding linear array. Unlike the random selection or projection-based approaches discussed earlier, these matrix completion and interpolation methods directly acquire sparse measurements. However, these techniques still complete the missing measurements before estimating target parameters, leading to increased computational complexity at the receiver. Contrarily, our approach estimates target parameters directly from the support of the recovered sparse signal, without requiring complete measurements. This enables not only reduced computational complexity and hardware requirements but also achieves superior resolution compared to conventional methods.
@@
Alternatively, spatial CS enables direct parameter recovery from SLAs[32, 33]. For instance, [32] focused on pulsed-MIMO radar and estimated AOAs for a specific range-velocity bin using different CS recovery algorithms. Additionally, [34] incorporated velocity estimation but not range. In [33], velocities and angles were estimated using measurements from a small number of randomly positioned transmitters and receivers on a circular disc. Besides spatial compression, CS techniques have also been applied in radar systems for interference mitigation [35], spectrum sharing [36], joint radar-communication systems[37], and multi-target shadowing effect mitigation[38]. Furthermore, [39] discussed the advantages and challenges of applying CS in radar, including clutter cancellation. However, these earlier studies focused primarily on pulsed-wave radars and were limited to estimating at most two target parameters. Conversely, our work develops CS-based techniques for estimating range, velocity, and AOA in MIMO-FMCW radars. To this end, we introduce a joint 2D-CS framework for Doppler-angle recovery, which has not been previously explored in the literature.
@@ I-B Our contributions
Preliminary results of this work appeared in our conference publication [1], where we addressed range and angle estimation only, without any theoretical guarantees. In this work, we consider a MIMO-FMCW radar with a random SLA that transmits only a subset of chirps per CPI and present a joint Doppler-angle estimation framework. The reduced number of antenna elements and chirps correspond to spatial and slow-time CS, respectively. To the best of our knowledge, the use of randomly transmitted sparse chirps, their integration with random SLAs for joint Doppler–angle estimation in MIMO-FMCW radars, and the associated recovery guarantees have not been investigated previously. These aspects constitute the key novel contributions of our work, as follows:
1) Range estimation: Prior works on CS-based radars have mainly focused on pulsed-wave radars and/or recovering target angles and velocities within a specific range bin. In this paper, we first present a separable mixture model for the IF signal, applicable to both full and sparse measurements. We propose two methods for range estimation: (a) DFT-focusing followed by binary integration[10], and (b) range-orthogonal matching pursuit (Range-OMP). In contrast to the coherently-integrated chirps in the conventional approaches, binary integration enhances detection performance at low SNRs, while the focusing operation[27] concentrates all the target returns from the same range in a single DFT-bin irrespective of their velocities and AOAs. However, in both conventional and proposed DFT-based range estimation, resolution is tied to the DFT-defined range bins such that achieving finer resolution requires increased bandwidth and results in higher computational costs. To this end, we propose Range-OMP, which uses a user-defined grid and a greedy OMP algorithm to deliver higher resolution with lower computational complexity while relying on measurements from just a single chirp and array channel; see further Remarks 2-4. Unlike standard OMP, which reconstructs sparse signals from compressed measurements, our Range-OMP leverages full fast-time radar measurements obtained through uniform sampling of the IF signal.
2) Joint Doppler-angle estimation: We jointly estimate target velocities and AOAs within each detected range bin using both vectorized and joint 2D-CS techniques. In conventional radar systems, achieving finer velocity and angular resolution typically demands transmitting a large number of chirps per CPI and deploying wide aperture arrays with many physical antenna elements—both of which significantly increase computational burden. Although super-resolution methods like MUSIC can enhance resolution without increasing measurements, they generally incur higher computational costs compared to simpler DFT-based approaches. To overcome these challenges, we adopt sparse random chirps and random SLAs, thereby reducing both slow-time and spatial measurements, and develop a joint CS-based velocity-angle estimation framework. Despite this reduction, our methods achieve performance on par with full-measurement systems, aided by user-defined grids that enable high-resolution estimation of velocity and AOA. Furthermore, the reduced measurements lower computational complexity and allow the radar to scan multiple angular sectors within a single CPI. We further summarize these advantages over conventional methods in Table I.
@@
Throughout the paper, we reserve boldface lowercase and uppercase letters for vectors (column vectors) and matrices, respectively, and $\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$ (or simply $a_{i_{1}\leq i\leq i_{2}}$) denotes a set of elements indexed by an integer $i$. The notations $\mathbf{A}]_{:,i}$, $\mathbf{A}]_{i,:}$ and $\mathbf{A}]_{i,j}$ denote the $i$-th column, $i$-th row and $(i,j)$-th element of matrix $\mathbf{A}$, respectively. The transpose/ Hermitian/conjugate operation is $(\cdot)^{T/H/*}$, expectation/probability is $\mathbb{E}/\mathbb{P}[\cdot]$, the outer product is $\otimes$ and the complement of a set is $(\cdot)^{c}$. The $l_{2/1/0}$ norm of a vector is $\|\cdot\|_{2/1/0}$. The notations supp, diag, and vec denote support, diagonal matrix, and vectorization, respectively. Also, $\mathbf{I}_{n}$ and $\mathbf{0}$ denote a $n\times n$’ identity matrix and an all-zero matrix, respectively. We represent the circular-normal and real-valued normal distributions as $\mathcal{CN}(\bm{\mu},\mathbf{Q})$ and $\mathcal{N}(\bm{\mu},\mathbf{Q})$ (with mean $\bm{\mu}$ and covariance matrix $\mathbf{Q}$) while $\mathcal{U}[a,b]/\mathcal{U}\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$ represents a uniform/ discrete-uniform distribution over interval [a,b][a,b]/ set $\{a_{i}\}_{i_{1}\leq i\leq i_{2}}$.
@@ II System Model
The transmitters transmit orthogonal LFM chirps of carrier frequency $f_{c}$, chirp rate $\gamma$, and chirp duration $T_{c}$. The wavelength is $\lambda=c/f_{c}$, where $c$ represents the speed of light. For simplicity, we consider time-domain orthogonality, i.e., the transmitters transmit the same signal with relative time shifts. Hence, the received signal components corresponding to various transmitters are readily separated at different receivers. Alternatively, orthogonal waveforms for MIMO-FMCW radars have been designed using beat frequency and chirp rate offsets [40, 41]. In [26], difference co-chirps are proposed for high-accuracy and low complexity range-Doppler estimation in FMCW radars. We consider a CPI of duration $T_{p}=P_{max}T_{c}$ corresponding to $P_{max}$ chirps. The standard radar system with full (slow-time) measurements transmits for the entire CPI. In this work, the radar transmits only $P$ out of $P_{max}$ chirps ($P<P_{max}$) in a specific angular sector of interest. We denote the sparse chirp set by $\mathcal{P}_{s}=\{\zeta_{p}\}_{1\leq p\leq P}$ where $\zeta_{p}$ are randomly drawn distinct integers from $\{0,1,\ldots,P_{max}-1\}$ and distributed as $\mathcal{P}_{p}$. On the other hand, in the standard radar, the set $\{\zeta_{p}\}=\{0,1,\ldots,P_{max}-1\}$ and all chirps are transmitted in the same angular sector. In our sparse radar setup, the time corresponding to $\mathcal{P}_{s}^{c}$, i.e., the non-transmitted chirps, can be utilized in other angular sectors, as depicted in Fig. 1b. Throughout the paper, we consider the target’s parameter estimation in a single sector only. The received signals from other sectors can be trivially processed in the same manner.
@@
$s(t)=\exp{\left(j2\pi\left(f_{c}t+\frac{\gamma}{2}t^{2}\right)\right)},\;\;0\leq t\leq T_{c}.$
@@
where $a_{k}$ is the complex amplitude proportional to the $k$-th target’s radar cross-section (RCS) and $\tau^{k}_{n,m,p}$ is the total delay in the $k$-th target’s return. The delay $\tau^{k}_{n,m,p}$ consists of range delay $\tau^{R}_{k}$, Doppler delay $\tau^{D}_{p,k}$ and angular delay $\tau^{\theta}_{n,m,k}$ as
@@
$\;\;\times\exp{\left(-j\pi\gamma(\tau^{k}_{n,m,p})^{2}\right)}\exp{\left(j2\pi f_{c}\tau^{k}_{n,m,p}\right)},$
@@
$\times\underbrace{\exp{\left(-j\pi\gamma(\tau^{k}_{n,m,p})^{2}\right)}}_{\textrm{Term-II}}\underbrace{\exp{\left(j2\pi f_{c}\tau^{k}_{n,m,p}\right)}}_{\textrm{Term-III}}.$ (2)
We, henceforth, deal with only discrete-time signals and use $t$ to denote the discrete-time index with $0\leq t\leq N-1$ where $N=\lceil f_{s}T_{c}\rceil$. In a MIMO-FMCW radar with $N_{T}$ transmitters, $N_{R}$ receivers and $P$ chirps, we obtain $N_{T}N_{R}\times P$’ sampled measurements $\{y_{n,m,p}[t]\}_{1\leq n\leq N_{T},1\leq m\leq N_{R},1\leq p\leq P}$.
@@
Term-I: For practical FMCW radars with narrow-band assumption, we have $\tau^{R}_{k}\gg\tau^{D}_{p,k}$ and $\tau^{R}_{k}\gg\tau^{\theta}_{n,m,k}$ for all targets such that $\gamma\tau^{k}_{n,m,p}/f_{s}\approx\gamma\tau^{R}_{k}/f_{s}$. This can also be verified numerically for a target with the radar’s parameters provided in Table III. We define the normalized beat frequency (due to range) as $\Omega_{R}^{k}\doteq\gamma\tau^{R}_{k}/f_{s}$ and approximate Term-I as $\exp{(j2\pi\Omega_{R}^{k}t)}$. This term is further examined in Section III-B using DFT-based focusing or OMP-recovery methods to obtain the range estimates.
Term-II: Again, using $\tau^{R}_{k}\gg\tau^{D}_{p,k}+\tau^{\theta}_{n,m,k}$, we have $(\tau^{k}_{n,m,p})^{2}\approx(\tau^{R}_{k})^{2}$ [42]. Hence, Term-II becomes $\exp{(-j\pi\gamma)(\tau^{R}_{k})^{2}}$, a constant for each target independent of chirp and antenna indices. In the 3D-mixture model (3), we include this term in $\widetilde{a}_{k}$, which is not the focus of the parameter estimation problem addressed in this paper.
@@
$\;\;\times\exp{\left(j2\pi\frac{A\sin(\theta_{k})}{2\lambda}(\alpha_{n}+\beta_{m})\right)}.$
@@
With these approximations and including the circular-normal noise term $w_{n,m,p}[t]$ to represent the interference and noises present in the radar measurements, (2) simplifies to
@@
$\;\;\times\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}+w_{n,m,p}[t],$ (3)
where $\widetilde{a}_{k}=a_{k}exp(j\pi\gamma(\tau^{R}_{k})^{2})exp(-j2\pi f_{c}\tau^{R}_{k})$. Here, (3) consists of a separable mixture of three different complex exponentials such that the target’s range, velocity, and AOA estimation problem is equivalent to 3D frequency estimation.
Remark 1 (Generalized model).
@@
Limited measurements: In a typical radar system estimating range, velocity, and AOA, three distinct types of measurements are used: (a) fast-time samples, i.e., the sampled IF signal, used for range estimation, (b) slow-time samples obtained from multiple chirps within a CPI for velocity estimation, and (c) spatial samples collected across array channels for AOA estimation. In our work, we adopt the conventional fast-time sampling at frequency $f_{s}$, allowing for range estimation through DFT-based focusing method in Section III-B.1. Both classical-DFT and our DFT-focusing rely on DFT-defined range bins, which influence both resolution and computational cost. In contrast, Section III-B.2 presents the Range-OMP algorithm, which utilizes the same fast-time data to achieve finer range resolution with reduced complexity. However, unlike full-measurement systems, we limit the number of slow-time and spatial domain samples. Specifically, only $P<P_{max}$ chirps are transmitted per CPI, and our SLA, illustrated in Fig. 1a, comprises of $N_{T}<N_{T,f}$ transmitters and $N_{R}<N_{R,f}$ receivers placed across apertures $A_{T}$ and $A_{R}$, respectively. Here, $N_{T,f}$ and $N_{R,f}$ denote the total transmitters and receivers in a full ULA. Traditional approaches like classical-DFT and MUSIC leverage complete measurements from all $P_{max}$ chirps and the full $N_{T,f}\times N_{R,f}$ array channels, and hence, improved resolution comes at the cost of increased measurement and processing overhead. To address this, we employ CS techniques in Section III-C to jointly estimate velocity and AOA with enhanced resolution and significantly lower computational load, despite relying on fewer chirps and antenna elements. As demonstrated in Section V, our proposed methods achieve superior estimation accuracy over classical-DFT and MUSIC, even when using only half the number of chirps and array elements compared to the full-measurement radar configuration.